find the slope-intercept form for the line satisfying x-intercept 2, y-intercept 2/3. The slope-intercept form for the line is?
Answer by fcabanski(1391)

y=mx+b is the slope intercept form. m is the slope. b is the y intercept.


The problem does not give the slope directly. But it shows two points on the line. The y-intercept is the point where x=0. So that point is (0,2/3). The x-intercept is the point where y=0. So that point is (2,0). Use those points to find the slope change in y / change in x.


(2/3 - 0) / (0-2) = 2/3 / -2 = - 2/6 = -1/3


y = -1/3 x + 2/3

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)